OpenStudy (anonymous):

The coordinates of the vertices of a square are A (-2, 5), B(3, 6), C(4, 1), and D(-1, 0). Find the coordinates of the midpoints of its sides and the midpoint of its diagonals. (Use fraction format.) Midpoint of side: AB = (, ) Midpoint of side: BC = (, ) Midpoint of side: CD = (, ) Midpoint of side: DA = (, ) Midpoint of diagonals: (, )

OpenStudy (anonymous):

AB=(1/2,11/2) BC=(7/2,7/2) CD=(3/2,1/2) DA=(-3/2,5/2) diagonal =(1,3)